OpenStudy (anonymous):

how can you express 499 to find the product 499x5 using mental math? explain.

OpenStudy (anonymous):

Another way to look at the same problem is perhaps this. We will try an example of 398x3. This is the same as 3(398). 3(398) = 3(400-2) 3(400-2) = 1200-6 1200-6 is 1194 It is basically taking what you have and changing it to numbers you can mentally visualize better and working with those.

OpenStudy (anonymous):

@coolkid17 Let me know if this helps at all. I can help break it down if you want.